Find the missing term
5,6,9,15,?,40

Answer by Theo(13342)

it looks to me like the interval between each of the number is the sum of the terms up to that number.

term 1 is equal to 5.
term 2 is equal to 5  + (1)                 = 5 + 1   = 6
term 3 is equal to 6  + (1 + 2)             = 6 + 3   = 9
term 4 is equal to 9  + (1 + 2 + 3)         = 9 + 6   = 15
term 5 is equal to 15 + (1 + 2 + 3 + 4)     = 15 + 10 = 25
term 6 is equal to 25 + (1 + 2 + 3 + 4 + 5) = 25 + 15 = 40


each succeeding term is the sum of the value of the previous term plus the sum of the value of the term numbers from 1 to the term number of the previous term.

for example, the value of term number 3 is the sum of the value of the previous term, which is 6, plus the sum of the value of the term numbers from 1 to the term number of the previous term, which is 2.
add the term numbers from 1 to 2 and you get 3.
add 3 to 6 and you get the value of term number 3 which is 9.

it's convoluted, but the logic seems to work so i'll go with it.